OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is perpendicular to the line y=3/5x-1 and passes through the point (-9, 4). (the 3 is on top of the 5)

OpenStudy (triciaal):

equation of a line in slope intercept form y = mx + b perpendicular lines the product of the slopes is -1 given the slope of the line is 3/5 then the slope of the perpendicular = -1/(3/5)= -5/3 the line passes through the point (x. y) any point on the perpendicular line and (-9, 4) known point y-4= -5/3(x--9) y -4+ 4 =-5/3(x) + 45/3 + 4 y = -5/3(x) + 19

OpenStudy (anonymous):

thank you!

OpenStudy (triciaal):

wait I have an error

OpenStudy (triciaal):

+45/3 should be -45/3